The length of stay of patients at a certain hospital in Cavite is 3 days with a standard deviation of 1.5 days. If a patient is selected at random, what is the probability that he/she will stay in the hospital between 3 and 7 days?


1
Expert's answer
2022-03-24T11:27:08-0400

We have a normal distribution, "\mu=3, \sigma=1.5."

Let's convert it to the standard normal distribution, "z=\cfrac{x-\mu}{\sigma};"

"z_1=\cfrac{3-3}{1.5}=0, \\z_2=\cfrac{7-3}{1.5}=2.67,"


"P(3<X<7)=P(0<Z<2.67)=\\=P(Z<2.67)-P(Z<0)="

"=0.9962-0.5=0.4962" ​(from z-table)